OpenStudy (anonymous):

solve the system of equations {3x-y=6 { 6x+y=21 explain the steps on how to do this please

OpenStudy (uri):

Easy

OpenStudy (uri):

Add both the equations

OpenStudy (anonymous):

x=3

OpenStudy (uri):

-y +y will be cancelled.

OpenStudy (uri):

3x-y=6 6x+y=21 ====== 9x=27 x=27/9 x=3

OpenStudy (mathstudent55):

There are different methods of solving systems of equations. Since you don't mention a specific method, I suggest addition. Add the two equations. The y variables will add to zero, leaving an equation with only the x variable. Then solve the equation for x.

OpenStudy (uri):

Now put the value of x in eq no.2 6x+y=21 6(3)+y=21 18+y=21 y=21-18 y=3 (3,3)

OpenStudy (anonymous):

thank you so much that really helped

OpenStudy (uri):

yw